OpenStudy (anonymous):

What is cos q if sec q = 2?

Nnesha (nnesha):

cos and sec are reciprocal of each other

Nnesha (nnesha):

\[\rm sec = \frac{ 1 }{ \cos }\]\[\cos =\frac{ 1 }{ \sec }\] 2 is same as 2/1 so if sec =2/1 cos =what ?

OpenStudy (anonymous):

1/2

Nnesha (nnesha):

yes right